what is the no. of wave in fourth orbit-
a.4  b.5  c.0  d.1

Grade:

1 Answers

Arun
5 years ago

No.of waves = circumference of electron orbit / wave length

                          =2πr /λ
                          λ=h /mv
               No.of waves =2π/h *(mvr)
         The angular momentum in 4th orbit mvr = nh/2π
                   = 4h/2π
            Substitute the mvr value
                 No.of waves=2π/h* 4h/2π =4